OpenStudy (anonymous):

can someone plz help! Find all solutions of the equation in the interval [0,2pi) 2cos(x+pi)+2cos(x-pi)=0 the answers r supposed to be pi over 2 and 3pi over 2 i dont get it!

myininaya (myininaya):

\[2[\cos(x)\cos(\pi)-\sin(x) \sin(\pi)]+2[\cos(x)\cos(\pi)+\sin(x) \sin(\pi)]\] \[=2[-\cos(x)]+2[-\cos(x)]=-2\cos(x)-2\cos(x)=-4\cos(x)=0\]

myininaya (myininaya):

\[=> \cos(x)=0\] I will leave the rest to you :)

OpenStudy (anonymous):

thnx a million!:))))))))